# GVE-Leiden: Fast Leiden Algorithm for Community Detection in Shared Memory Setting

### v7

@@ 1. Introduction
In this report, we present our parallel multicore implementation of the Louvain algorithm [^1] https://github.com/puzzlef/leiden-communities-openmp. It incorporates several optimizations, including parallel prefix sums, preallocated Compressed Sparse Row (CSR) data structures for community vertex identification and super-vertex graph storage during aggregation, fast collision-free per-thread hash tables for the local-moving and aggregation phases, and prevention of unnecessary aggregations — enabling it to run significantly faster than existing implementations. Additionally, we employ a greedy refinement phase where vertices optimize for delta-modularity within their community bounds, yielding improved performance and quality compared to a randomized approach. Furthermore, we utilize established techniques such as asynchronous computation, OpenMP’s dynamic loop schedule, threshold-scaling optimization, and vertex pruning. To the best of our knowledge, our implementation is the most efficient implementation of Leiden algorithm on multicore CPUs to date. We conduct comprehensive comparisons with other state-of-the-art implementations, including multi-core, GPU-based implementations, multi-node implementations, detailed in Table 1. Both direct and indirect comparisons are provided, with details given in Sections 5.2 and A.2, respectively.
@@
Original Leiden (Traag et al., 2019) Sequential $\times2$
igraph Leiden (Csardi et al., 2006) Sequential 50×50$\times$50 ×
NetworKit Leiden (Staudt et al., 2016) Parallel 20×20$\times$20 ×
@@
ParLeiden-S (Hu et al., [n. d.]) Parallel $22\times^{*}$
ParLeiden-D (8 nodes) (Hu et al., [n. d.]) Multi-node $18\times^{*}$

### v8

@@ 1. Introduction
In this report, we present our parallel multicore implementation of the Leiden algorithm [^1] https://github.com/puzzlef/leiden-communities-openmp. It incorporates several optimizations, including parallel prefix sums, preallocated Compressed Sparse Row (CSR) data structures for community vertex identification and super-vertex graph storage during aggregation, fast collision-free per-thread hash tables for the local-moving and aggregation phases, and prevention of unnecessary aggregations — enabling it to run significantly faster than existing implementations. Additionally, we employ a greedy refinement phase where vertices optimize for delta-modularity within their community bounds, yielding improved performance and quality compared to a randomized approach. Furthermore, we utilize established techniques such as asynchronous computation, OpenMP’s dynamic loop schedule, threshold-scaling optimization, and vertex pruning. To the best of our knowledge, our implementation is the most efficient implementation of Leiden algorithm on multicore CPUs to date. We conduct comprehensive comparisons with other state-of-the-art implementations, including multi-core, GPU-based implementations, multi-node implementations, detailed in Table 1. Both direct and indirect comparisons are provided, with details given in Sections 5.2 and A.2, respectively.
@@
Original Leiden (Traag et al., 2019) Sequential 436×436$\times$436 ×
igraph Leiden (Csardi et al., 2006) Sequential 104×104$\times$104 ×
NetworKit Leiden (Staudt et al., 2016) Parallel 8.2×8.2$\times$8.2 ×
@@
ParLeiden-S (Hu et al., [n. d.]) Parallel $18\times^{*}$
ParLeiden-D (8 nodes) (Hu et al., [n. d.]) Multi-node $22\times^{*}$
@@ 2. Related work
However, most existing works only focus on optimizing the local-moving phase of Leiden algorithm, are not sufficiently optimized, and are either sequential or lack effective parallelization. For instance, the original Leiden implementation (Traag et al., 2019), uses a full Graph data structure for each of the collapsed graphs without reuse and makes a complete copy of the communities when collapsing the graph. Further, it tracks the community membership of vertices in each dendrogram layer, adding to the computational overhead. Additionally, it employs a queue and a flag vector to monitor the nodes to be moved in each iteration, with the processing order randomly shuffled. A new degree vector is allocated in each iteration, and community memberships are renumbered after each local-moving and refinement phase. This implementation also conducts several book-keeping activities when moving a node to another community, such as maintaining community sets, removing empty communities, and updating the total edge weight from communities. Finally, it also keeps track of nodes with fixed community memberships using three vectors, a process repeated in each local-moving phase.
The igraph Leiden implementation (Csardi et al., 2006) follows a comparable approach but with some differences that streamline its operations slightly. Like the original Leiden, it uses a full Graph data structure for each collapsed graph without reuse. igraph Leiden also uses a queue and a flag vector to manage nodes during the local-moving phase, with the processing order randomly shuffled as well. In addition, it also reindexes the community membership of vertices after the local-moving phase, and the refinement phase also involves a random node shuffling. However, unlike the original Leiden, igraph only maintains the size of each community rather than the individual nodes, simplifying the process. Unlike the original Leiden, it does not track edge weights from communities but instead keeps track of total edge weights, which is easier to manage.
The NetworKit Leiden implementation (Staudt et al., 2016) introduces parallel processing but retains some of the complexity seen in the above implementations. It uses a queue and a flag vector to track vertices for processing, incorporating a mutex and a condition variable for parallel access and updates, with node shuffling done on a single thread. Synchronization between threads is performed to manage available work, while employing locks. During the aggregation phase, not all operations are parallelized, such as the prefix sum calculation. Community membership of vertices is tracked and flattened at the end, while node mapping is stored in a std::map, and fine to coarse mapping is sequentially stored. These issues contribute to the computational expense of NetworKit Leiden, despite the benefits of parallel processing.